# Leveraging MTD to Mitigate Poisoning Attacks in Decentralized FL with Non-IID Data

### v1

@@ Abstract
Decentralized Federated Learning (DFL), a paradigm for managing big data in a privacy-preserved manner, is still vulnerable to poisoning attacks where malicious clients tamper with data or models. Current defense methods often assume Independently and Identically Distributed (IID) data, which is unrealistic in real-world applications. In non-IID contexts, existing defensive strategies face challenges in distinguishing between models that have been compromised and those that have been trained on heterogeneous data distributions, leading to diminished efficacy. In response, this paper proposes a framework that employs the Moving Target Defense (MTD) approach to bolster the robustness of DFL models. By continuously modifying the attack surface of the DFL system, this framework aims to mitigate poisoning attacks effectively. The proposed MTD framework includes both proactive and reactive modes, utilizing a reputation system that combines metrics of model similarity and loss, alongside various defensive techniques. Comprehensive experimental evaluations indicate that the MTD-based mechanism significantly mitigates a range of poisoning attack types across multiple datasets with different topologies.
@@ I Introduction
Movement Target Defense (MTD) is an emerging cybersecurity paradigm for mitigating various attacks [7]. By dynamically altering the attack surface, either actively or passively, MTD increases the complexity and cost of successfully executing a cyberattack [8]. Thus, there is a need to apply MTD to enhance the security and robustness of DFL models, particularly in defending against poisoning attacks.
This work proposed an MTD-based approach to dynamically modify the attack surface of DFL systems, thereby improving their resilience and security. The main contributions of this paper are: (1) Designing an MTD-based approach to mitigate poisoning attacks in DFL. This approach utilizes dual-reputation scores based on model similarity and loss, adopts proactive and reactive modes, and incorporates dynamic topology and aggregation algorithms. (2) Employing the DBSCAN algorithm to adjust the triggering threshold for MTD dynamically. (3) Implementing a prototype of the proposed MTD method on a real DFL platform [9]. (4) Conducting extensive experiments to compare the proposed algorithm with state-of-the-art methods, including scenarios with fully connected, ring, and star topologies, as well as untargeted poisoning and backdoor attacks on MNIST, FashionMNIST, and CIFAR-10 datasets. The results demonstrate that the proposed MTD-based algorithm not only effectively mitigates poisoning attacks, outperforming existing methods, but also improves model performance on non-IID data.
The remainder of this paper is organized as follows: Section II provides a summary and review of research utilizing MTD to address AI security challenges. Section III outlines the problems this paper aims to solve and presents the threat model. Section IV details the design of our proposed MTD-based approach. Subsequently, Section V evaluates and compares the effectiveness of our method with state-of-the-art approaches in mitigating poisoning attacks. Section VI discusses the experimental results, and Section VII concludes the paper by summarizing key contributions and highlighting future research opportunities.
@@ II Background and Related Work
Robust aggregation functions are the most commonly adopted defense approaches in FL [4] against poisoning attacks. Krum [5] identifies the update that is nearest to the average of the remaining updates, deliberately excluding any outliers from consideration. Median [18] employs the median value of the updates, which exhibits reduced sensitivity to outlier effects. TrimmedMean [18] calculates the updates’ mean after excluding a predetermined proportion of extreme values. Nonetheless, these defenses predominantly utilize static strategies typically formulated for the conventional FL paradigm rather than for DFL. These static defenses are ineffective in addressing poisoning attacks targeting DFL systems, especially in non-IID environments [4].
@@
TABLE I provides an overview of the existing literature on applying MTD strategies in the context of Artificial Intelligence (AI) Security to address diverse cyber threats. In the traditional AI field, MTD has been employed to mitigate adversarial samples and poisoning attacks. [10] proposes an MTD approach to enhance Deep Neural Network (DNN) robustness by dynamically changing model parameters during inference. This makes it harder for attackers to craft compelling adversarial examples, improving resilience against white and black-box attacks. [11] introduces Morphence, an MTD strategy generating diverse student models from a base model with random weight perturbations. Randomly selecting a student model for each inference reduces the adversarial attack success rate while maintaining accuracy. [12] proposes EI-MTD, a dynamic defense mechanism for edge computing. It uses a three-stage framework to address challenges like limited resources and transferability, aiming to prevent adversaries from finding substitute models and reducing attack transferability. [13] presents MT-MTD, an MTD framework protecting DNNs against backdoor attacks in edge computing. It divides training data, has an attacker train each dimension, and the defender randomly selects dimensions for consensus, increasing attack cost and maintaining accuracy. [14] introduces MTDeep, enhancing neural network robustness using MTD by dynamically changing model architecture, making it harder for attackers to craft adversarial attacks.
@@
However, there is limited research on utilizing MTD methodology to address the security challenges of FL systems. [15] enhances FL security with MTD by introducing an augmented dual-shuffle mechanism. This dynamically changes the attack surface by shuffling client selection and model aggregation, improving system security against various poisoning attacks. [16] introduces a proactive MTD mechanism that continuously reconfigures the network topology to enhance DFL security. [17] proposes Voyager, which employs a reactive MTD-based aggregation protocol to bolster the resilience of DFL against poisoning attacks. This is achieved by dynamically manipulating the connectivity of the network topology. However, its efficacy remains to be validated in non-IID scenarios.
@@
III Problem Setup and Threat Model
@@ III-A Poisoning Attacks in non-IID Environments
III-B Threat Model
The threat model is essential for systematically analyzing and delineating the research problem. This paper examines the poisoning attack on DFL through the lens of the attacker’s objectives, knowledge, and capabilities.
@@ IV-B MTD Module
An MTD module can be characterized by a triplet ¡C, M, T¿, where C signifies the configuration space within which the MTD module is capable of moving, M indicates the strategy employed for the action, and T represents the timing function that determines the intervals at which actions are executed [20]. In essence, the design of an MTD module necessitates addressing three fundamental questions: (1) What to move; (2) How to move; and (3) When to move [21].
@@ VI Discussion
Extensive experiments have demonstrated that the MTD-based approach proposed in this paper can effectively mitigate poisoning attacks in non-IID environments, particularly for untargeted attacks. Therefore, the algorithm presented herein emerges as a dependable defense mechanism. Nevertheless, despite this approach demonstrating superiority over existing state-of-the-art algorithms, there is potential for enhancement in addressing targeted attacks, such as backdoor attacks.
This work analyzes the distribution of reputation scores generated by benign nodes to investigate the reasons behind the suboptimal performance of the MTD-based approach in backdoor attacks. It is conducted under conditions where the PNR is set at 10% and within a fully connected topology. The comparison is performed across three datasets, focusing on model poisoning and backdoor attack scenarios.
@@
It is essential to highlight that the algorithm presented in this study enhances the model performance of DFL in non-IID environments, especially when using the dynamic topology strategy. The average F1-Score achieved by the proposed approach reaches 0.95 for the MNIST dataset, 0.83 for FashionMNIST, and 0.70 for Cifar10 when using the reactive dynamic topology strategy, as illustrated in Figure4 and 5. These results indicate that the algorithm not only mitigates the impacts of poisoning attacks but also enhances model performance in non-IID settings. The effectiveness of the dynamic topology lies in its tendency to facilitate connections and aggregations among nodes that exhibit high similarity, thereby diminishing the impact of non-IID data.
@@ VII Conclusion and Future Work
This work presents an MTD-based approach to mitigate poisoning attacks in DFL, particularly in non-IID environments. The approach employs a novel dual-reputation score mechanism that evaluates both model similarity and loss. It incorporates proactive and reactive MTD modes coupled with dynamic topology and aggregation algorithm strategies. By utilizing the DBSCAN clustering algorithm, the proposed MTD strategy overcomes the limitations of traditional defense mechanisms that often rely on fixed thresholds. Instead, it dynamically adjusts thresholds and triggers based on the current state of the federation. Furthermore, the dynamic topology strategy promotes connections among nodes that are closely aligned with the local model, enhancing performance in non-IID scenarios. The effectiveness of this MTD-based approach is demonstrated through experiments on MNIST, FashionMNIST, and CIFAR-10 datasets, employing various initial topologies (fully connected, ring, and star) and comparing against label flipping, model poisoning, and backdoor attacks. The results show that the proposed approach effectively mitigates a range of untargeted attacks and outperforms state-of-the-art aggregation algorithms in defending against backdoor attacks. Consequently, this MTD-based approach represents a robust and reliable security framework for non-IID environments.
Future research will investigate integrating additional MTD strategies to comprehensively address the diverse landscape of poisoning attacks and enhance resilience against optimized and dynamic attack methodologies. Further optimization of the MTD-based approach, particularly in the context of backdoor attacks, will also be explored. Additionally, the research scope will be extended to related domains, including the application of the MTD paradigm to counter various other types of attacks targeting DFL systems, such as Membership Inference Attacks.

### v2

@@ Abstract
Decentralized Federated Learning (DFL), a paradigm for managing big data in a privacy-preserving and distributed manner, is still vulnerable to poisoning attacks where malicious clients tamper data or models. Current defense methods often assume Independently and Identically Distributed (IID) data across participants, which is unrealistic in real-world applications. In more realistic non-IID contexts, existing defensive strategies face challenges when distinguishing between models that have been compromised and those that have been trained on heterogeneous data distributions (non-IID), leading to diminished efficacy. In response, this paper proposes a framework that employs the Moving Target Defense (MTD) approach to bolster the robustness of DFL models. By continuously modifying the attack surface of the DFL system, the framework aims to mitigate poisoning attacks effectively. The proposed solution includes both proactive and reactive modes, utilizing a reputation system that combines metrics of model similarity and loss, alongside various defensive techniques. Comprehensive experimental evaluations indicate that the MTD-based mechanism significantly mitigates a range of poisoning attack types across multiple datasets with different federation topologies.
@@ I Introduction
Movement Target Defense (MTD) is an emerging cybersecurity paradigm for mitigating various attacks [7]. By dynamically altering the attack surface, either actively or passively, MTD increases the complexity and cost of successfully executing a cyberattack [8]. A limited number of studies have endeavored to employ MTD to mitigate security concerns within DFL [9, 10]. However, their performance is diminished in non-IID environments. Thus, there is a need to apply MTD to enhance the security and robustness of DFL models, particularly in defending against poisoning attacks.
This work proposes an MTD-based approach to dynamically modify the attack surface of DFL systems and improve their resilience and security. Therefore, the main contributions of this paper are: (1) Designing an MTD-based approach to mitigate poisoning attacks in DFL. This approach utilizes a dual-reputation score system based on model similarity and loss, adopts proactive and reactive modes, and incorporates dynamic topology and aggregation strategies. (2) Employing the DBSCAN clustering ‘algorithm to adjust the triggering threshold for MTD dynamically. (3) Implementing a prototype of the proposed MTD method on a real DFL platform [11]. (4) Conducting extensive experiments to compare the proposed algorithm with state-of-the-art methods, including scenarios with fully connected, ring, and star topologies, as well as untargeted poisoning and backdoor attacks on MNIST, FashionMNIST, and CIFAR-10 datasets. The results demonstrate that the proposed MTD-based algorithm not only effectively mitigates poisoning attacks, outperforming existing methods, but also improves model performance on non-IID data.
The remainder of this paper is organized as follows: Section II provides a summary and review of research utilizing MTD to address AI security challenges. Section III outlines the problems this paper aims to solve. Section IV details the design of our proposed MTD-based approach. Subsequently, Section V evaluates and compares the effectiveness of our method with state-of-the-art approaches in mitigating poisoning attacks. Section VI discusses the experimental results, and Section VII concludes the paper by summarizing key contributions and highlighting future research opportunities.
@@ II Background and Related Work
Robust aggregation functions are the most commonly adopted defense approaches in FL against poisoning attacks [4]. Krum [5] identifies the update that is nearest to the average of the remaining updates, deliberately excluding outliers from consideration. Median [18] employs the median value of the updates, which exhibits reduced sensitivity to outlier effects. TrimmedMean [18] calculates the updates’ mean after excluding a predetermined proportion of extreme values. Nonetheless, these defenses predominantly utilize static strategies typically formulated for the conventional FL paradigm rather than for DFL. These static defenses are ineffective in addressing poisoning attacks targeting DFL systems, especially in non-IID environments [4].
@@
TABLE I provides an overview of the existing literature on applying MTD strategies in the context of Artificial Intelligence (AI) Security to address diverse cyber threats. In the traditional AI field, MTD has been employed to mitigate adversarial samples and poisoning attacks. [12] proposes an MTD approach to enhance Deep Neural Network (DNN) robustness by dynamically changing model parameters during inference. This makes it harder for attackers to craft compelling adversarial examples, improving resilience against white and black-box attacks. Morphence [13] is an MTD strategy generating diverse student models from a base model with random weight perturbations. Randomly selecting a student model for each inference reduces the adversarial attack success rate while maintaining accuracy. EI-MTD [14] is a dynamic defense mechanism for edge computing. It uses a three-stage framework to address challenges like limited resources and transferability, aiming to prevent adversaries from finding substitute models and reducing attack transferability. MT-MTD [15] presents an MTD framework protecting DNNs against backdoor attacks in edge computing. It divides training data, has an attacker train each dimension, and the defender randomly selects dimensions for consensus, increasing attack cost and maintaining accuracy. MTDeep [16] enhances neural network robustness using MTD by dynamically changing model architecture, making it harder for attackers to craft adversarial attacks.
@@
However, there is limited research on utilizing MTD methodology to address the security challenges of FL systems. [17] enhances FL security with MTD by introducing an augmented dual-shuffle mechanism. This dynamically changes the attack surface by shuffling client selection and model aggregation, improving system security against various poisoning attacks. [9] introduces a proactive MTD mechanism that continuously reconfigures the network topology to enhance DFL security. Voyager [10] employs a reactive MTD-based aggregation protocol to bolster the resilience of DFL against poisoning attacks. This is achieved by dynamically manipulating the connectivity of the network topology. However, its efficacy remains to be validated in non-IID scenarios.
@@
III Scenario and Adversarial Setup
This section analyzes the challenges of mitigating poisoning attacks, particularly highlighting the distinct behaviors these attacks exhibit in IID and non-IID data environments. Additionally, it examines the adversarial setup by considering the attacker’s objectives, knowledge, and capabilities. This analysis provides a foundation for understanding the specific problem and scenario addressed in this paper.
@@ III-A Poisoning Attacks in non-IID Environments
III-B Adversarial Setup
This section examines poisoning attacks on DFL through the lens of the attacker’s objectives, knowledge, and capabilities.
@@ IV-A Reputation System
The model similarity and loss of neighboring models are integrated into a two-dimensional vector, forming the reputation system’s metrics. The reputation score serves as the basis for judging the trustworthiness of neighboring models.
@@ IV-B MTD Module
An MTD module can be characterized by a triplet $<C,M,T>$, where $C$ signifies the configuration space within which the MTD module is capable of moving, $M$ indicates the strategy employed for the action, and $T$ represents the timing function that determines the intervals at which actions are executed [20]. In essence, the design of an MTD module necessitates addressing three fundamental questions: (1) What to move; (2) How to move; and (3) When to move [21].
@@ VI Discussion
Extensive experiments have demonstrated that the MTD-based approach proposed in this paper can effectively mitigate poisoning attacks in non-IID environments, particularly for untargeted attacks. Therefore, the algorithm presented herein emerges as a reliable defense mechanism. Nevertheless, despite this approach demonstrating superiority over existing state-of-the-art algorithms, there is potential for enhancement in addressing targeted attacks, such as backdoor attacks.
To investigate the reasons behind the sub-optimal performance of the MTD-based approach in backdoor attacks, this work analyzes the distribution of reputation scores generated by benign nodes. It is conducted under conditions where the PNR is set at 10% and within a fully connected topology. The comparison is performed across three datasets, focusing on model poisoning and backdoor attack scenarios.
@@
It is essential to highlight that the proposed solution enhances the model performance of DFL in non-IID environments, especially when using the dynamic topology strategy. The average F1-Score achieved by the proposed approach reaches 0.95 for the MNIST dataset, 0.83 for FashionMNIST, and 0.70 for Cifar10 when using the reactive dynamic topology strategy, as illustrated in Figure4 and 5. These results indicate that the algorithm not only mitigates the impacts of poisoning attacks but also enhances model performance in non-IID settings. The effectiveness of the dynamic topology lies in its tendency to facilitate connections and aggregations among nodes that exhibit high similarity, thereby diminishing the impact of non-IID data.
@@ VII Conclusion and Future Work
This work presents an MTD-based approach to mitigate poisoning attacks in DFL, particularly in non-IID environments. The approach employs a novel dual-reputation score mechanism that evaluates both model similarity and loss. It incorporates proactive and reactive MTD modes coupled with dynamic topology and aggregation algorithm strategies. By utilizing the DBSCAN clustering algorithm, the proposed MTD strategy overcomes the limitations of traditional defense mechanisms that often rely on fixed thresholds. Instead, it dynamically adjusts thresholds and triggers based on the current state of the federation. Furthermore, the dynamic topology strategy promotes connections among nodes that are closely aligned with the local model, enhancing performance in non-IID scenarios. The effectiveness of this MTD-based approach is demonstrated through experiments on MNIST, FashionMNIST, and CIFAR-10 datasets, employing various initial topologies (fully connected, ring, and star) and comparing against label flipping, model poisoning, and backdoor attacks. The results show that the proposed approach effectively mitigates a range of untargeted attacks and outperforms state-of-the-art aggregation algorithms in defending against backdoor attacks. Consequently, this MTD-based approach represents a robust and reliable security framework for non-IID data.
Future research will investigate integrating additional MTD strategies to comprehensively address the diverse landscape of poisoning attacks and enhance resilience against optimized and dynamic attack methodologies. Further optimization of the MTD-based approach, particularly in the context of backdoor attacks, will also be explored. Additionally, the research scope will be extended to related domains, including the application of the MTD paradigm to counter various other types of attacks targeting DFL, such as Membership Inference Attacks.